The sum of three consecutive even integers is 366. What is the largest of the three integers?
Answer by nyc_function(2741)

x = first even integer (smallest)

x + 2 = second even integer (middle)

x + 4 = third even integer (largest)

Here is your equation:

x + x + 2 + x + 4 = 366

Can you take it from here?